OpenStudy (anonymous):

Solve using undetermined coefficients y'' + y = t(1 + sin t) I got the t term to be At+B but the I'm kind of stuck on the tsint term.

OpenStudy (anonymous):

Is the related equation supposed to be t(Ct+D)(cos t + sint t)?

OpenStudy (eyust707):

You have to find the particular solution and the complementary solution and then use the thrm that says \(y(t) = y_{c} + y_{p}\) http://www.wolframalpha.com/input/?i=y%27%27+%2B+y+%3D+t%281+%2B+sin+t%29

OpenStudy (anonymous):

I know that I need to find the homogeneous solution and the particular solution. I'm stuck at finding the relationship between y'' + y = t sin (t) If it were just the sin term I could just use the e^t(A sin t + B cos t) (r=+/- i) relationship but I'm not sure about the t that's tacked on. Would it just be (Ct+D)(E sin t + F cos t)? Also, I don't actually need to evaluate the terms, just find the variable solution.

OpenStudy (experimentx):

i don't think you can solve that using the method of undetermined coefficients. use superposition, solve for t using undetermined coefficients. you can solve for the other using reduction of order or variation of parameters.

OpenStudy (anonymous):

Hmm I'm looking at some-what similar equation right now (y'' + y = 3 sin 2t + t cos 2t) and they found the solution using undetermined coefficients but the relation they used was y = t e^(rt) instead Acost+Bsint. Fixed a typo*

OpenStudy (experimentx):

find this y'' + y = t using method of undetermined coefficients. the particular solution will be t then find the particular solution of y'' + y = t sin(t)

OpenStudy (anonymous):

Yeah, that's what I did. I'm just having a hard time finding the relating equation to y'' + y = t sin(t)

OpenStudy (experimentx):

interesting .. seems i didn't know this http://en.wikipedia.org/wiki/Method_of_undetermined_coefficients#Examples

OpenStudy (anonymous):

Oh damn, I should of checked wiki... haha.

OpenStudy (experimentx):

seems that way ...

OpenStudy (anonymous):

apparently I don't need to differentiate y_2 and find the complete particular solution. That'll save me like 20 min of working through the alg portion of the problem haha.

OpenStudy (experimentx):

well ... this isn't quite nice method http://upload.wikimedia.org/math/7/0/a/70a575f1c539021bd893575c05ff365d.png

OpenStudy (experimentx):

seems both forms have the same solution type ... you can directly compare this result to t sin(t) and get the coefficients.

OpenStudy (anonymous):

I actually didn't need to go through the algebraic portion of the problem haha. The only part that was needed was this = t (Ct+D)sin t + t (Et+F) cos t. So the final solution for the problem is y = y1 + y2 = (At + B) + t[ (Ct+D)sin t + (Et+F) cos t ]

OpenStudy (experimentx):

you need to determine the values of A, B, C, D ... the second order equation cannot have more than two coefficients.

OpenStudy (anonymous):

I mean the problem actually only required the portion I typed out. The solution in the back of the book is actually the same except with A_0, A_1, B_0, B_1, etc. lol The second part of the question asked that we use a CAS to determine the coefficients. http://imageshack.us/a/img17/7984/asdasdasdwk.png

OpenStudy (experimentx):

well i guess then it is.